OpenStudy (anonymous):

Can someone check my answer please divide (8x^2-24x+15)/(2x-5) 4x-2+5/2x-5--- my answer

OpenStudy (anonymous):

@sweetburger do you mind checking this aswell?

hero (hero):

8x^2 - 24x + 15 = 8x^2 - 20x - 4x + 10 + 5 = 4x(2x - 5) - 2(2x - 5) + 5 = (2x - 5)(4x - 2) + 5

hero (hero):

Factoring is an alternative to long division that gets you the same result.

OpenStudy (anonymous):

thanks!